OpenStudy (anonymous):

i^-15 in standard form

OpenStudy (p0sitr0n):

i^-1?

OpenStudy (anonymous):

probably easiest to understand if you know \(i^{15}\) first do you know that one?

OpenStudy (jonnyvonny):

\[i^1=\sqrt{-1}\left| \right| i^2=\sqrt{-1}*\sqrt{-1}=1\left| \right| i^3=\sqrt{-1}*\sqrt{-1}*\sqrt{-1}=\sqrt{-1}\] The pattern continues. All negative does is makes "-1" become the denominator, which, essentially doesn't change anything, because 1/1, when you take the reciprocal, is still 1/1. Anyways, every pair you can cancel out because it results in 1, which when multiplied by anything, is still the multiplied value: 1*2=2....1*10x=10x ect. So yes:\[i^-=i^1=i\]

OpenStudy (anonymous):

hmm i would do this if i want \(i^{15}\) i think "4 goes in to 15 with a remainder of 3 and so \(i^{15}=i^3=-i\)

OpenStudy (anonymous):

by the way the pattern is not quite what was written above it is \[i^1=i\\ i^2=-1\\ i^3=-i\\ i^4=1\] and it repeats every 4

OpenStudy (anonymous):

this tells you \[i^{-15}=\frac{1}{-i}\] then \[\frac{1}{-i}\times \frac{i}{i}=i\]